### v1

# GCR: Geometry-Consistent Routing for Task-Agnostic Continual Anomaly Detection

Abstract
Feature-based anomaly detection is widely adopted in industrial inspection due to the strong representational power of large pre-trained vision encoders. While most existing methods focus on improving within-category anomaly scoring (e.g., via density modeling or nearest-neighbor retrieval), practical deployments increasingly require task-agnostic operation under continual category expansion, where the category identity is unknown at test time. In this setting, overall performance is often dominated by expert selection: routing an input to an appropriate normality model before any head-specific scoring is applied. However, routing rules that compare head-specific anomaly scores across independently constructed heads can be unreliable, since score distributions may differ substantially across categories in scale and tail behavior, leading to unstable decisions in multi-head continual evaluation.
We propose GCR, a lightweight mixture-of-experts framework that stabilizes task-agnostic continual anomaly detection via geometry-consistent routing. GCR routes each test image directly in a shared frozen patch-embedding space by minimizing an accumulated nearest-prototype distance to category-specific prototype banks, and then computes anomaly maps only within the routed expert using a standard prototype-based scoring rule. By explicitly separating cross-head decision making (routing) from within-head anomaly scoring, GCR avoids cross-head score comparability issues without requiring end-to-end representation learning.
Extensive experiments on MVTec AD and VisA demonstrate that geometry-consistent routing substantially improves routing stability and mitigates continual performance collapse, achieving near-zero forgetting while maintaining competitive detection and localization performance. Our results suggest that a large fraction of failures previously attributed to representation forgetting can instead be explained by decision-rule instability in cross-head routing, highlighting routing design as a key factor for robust task-agnostic continual anomaly detection. Code is publicly available at https://github.com/jw-chae/GCR
1 Introduction
Feature-based anomaly detection (AD) has become a widely adopted paradigm in industrial visual inspection, owing to its strong empirical performance and practical accessibility: a pre-trained vision encoder provides transferable patch representations, and anomalies are detected by measuring deviations from normal feature patterns without requiring task-specific supervision [bergmann2019mvtec, defard2021padim, roth2022towards]. A key implication of this paradigm is that, when the feature space is sufficiently informative, competitive detection and localization can be achieved without end-to-end representation learning, using simple scoring rules on top of frozen embeddings [defard2021padim, roth2022towards].
However, real industrial systems rarely operate in a single-category regime. A single production line often manufactures dozens of products whose visual patterns differ subtly, and test-time inference is frequently performed without access to category labels or reliable meta-information. This motivates task-agnostic anomaly detection, where a system must handle multiple normal manifolds while making decisions without category supervision at inference time [you2022unified, gao2024learning, he2024diffusion, he2024learning].
In addition, product portfolios evolve: new categories are introduced over time, and industrial inspection systems are expected to incorporate new normal patterns while maintaining performance on previously deployed categories. This motivates continual anomaly detection, where categories arrive sequentially and the system must avoid severe degradation on earlier categories after learning new ones [liu2024unsupervised]. A recurring empirical observation in such settings is that introducing new categories can dramatically reduce performance on past categories under task-agnostic evaluation, a phenomenon often described as catastrophic forgetting at the system level [liu2024unsupervised]. Recent approaches such as DFM and UCAD address this challenge by adopting multi-head designs and routing mechanisms, and further attempt to improve routing via additional representation learning or auxiliary objectives [wu2025dfm, liu2024unsupervised].
This raises a fundamental question: if the feature space is already sufficiently discriminative, is additional representation learning truly necessary for routing stability? In a multi-head system, once per-category experts are constructed, overall success under task-agnostic continual evaluation depends critically on expert selection—routing each test input to an appropriate head before applying any head-specific anomaly scoring [liu2024unsupervised]. Therefore, it is natural to ask whether the observed continual collapse primarily reflects insufficient representational capacity, or whether it can largely be attributed to decision instability in cross-head routing when head-wise scores are heterogeneous [you2022unified, liu2024unsupervised].
To probe this hypothesis, we conduct a diagnostic study using PatchCore-style retrieval on top of frozen patch embeddings from a pre-trained ViT backbone. We construct per-category prototype banks from normal training patches and perform task-agnostic inference by routing each test image through a shared geometric criterion in the frozen embedding space, followed by standard within-head anomaly scoring [roth2022towards]. Surprisingly, this routing strategy yields near-no performance degradation relative to category-aware evaluation in our setup, suggesting that the frozen feature space provides sufficient separability for both category discrimination and anomaly detection when routing is geometry-consistent [roth2022towards]. This observation, together with the strong sensitivity of continual performance to routing behavior reported in prior work, suggests that a substantial portion of the “catastrophic forgetting” observed in task-agnostic continual AD may be better understood as cross-head decision instability rather than representational degradation [liu2024unsupervised].
Motivated by this perspective, we propose GCR (Geometry-Consistent Routing) for task-agnostic continual anomaly detection. GCR adopts a multi-head Mixture-of-Experts view and decouples cross-head routing from within-head anomaly scoring: routing is performed using a shared geometric criterion in the frozen patch-embedding space, and anomaly maps are computed only within the selected expert. By focusing on routing design as a decision-rule problem, GCR improves routing stability under heterogeneous heads without requiring end-to-end representation learning [liu2024unsupervised, you2022unified].
Our contributions are threefold:
  • A decision-rule perspective on task-agnostic continual AD. We view task-agnostic continual anomaly detection as a multi-head expert selection problem and discuss how performance collapse under continual evaluation can be associated with cross-head decision instability in routing, in addition to representational factors [you2022unified, liu2024unsupervised].
  • GCR: geometry-consistent routing as a lightweight MoE design. We propose GCR, a lightweight Mixture-of-Experts framework that separates cross-head routing from within-head anomaly scoring. GCR routes inputs using a shared geometric criterion in a frozen patch-embedding space with per-category prototype banks, and evaluates anomaly scores only within the selected expert, avoiding direct cross-head comparisons of heterogeneous head scores [roth2022towards, liu2024unsupervised].
  • Evaluation with routing diagnostics. We evaluate GCR on MVTec AD and VisA under task-agnostic continual protocols and report routing diagnostics (e.g., routing accuracy and conditional performance) to help disentangle routing errors from within-head scoring behavior. The results indicate that geometry-consistent routing is associated with improved routing stability and reduced continual performance collapse while retaining competitive detection and localization performance [bergmann2019mvtec, liu2024unsupervised, wu2025dfm].
Figure 1: Overview of GCR. (a) Training Phase) GCR extracts multi-layer patch features from a frozen Vision Transformer and concatenates them to form patch-level representations. For each category, a prototype memory bank is constructed via coreset selection to compactly represent normal patterns, without explicit parameter learning or anomaly scoring during training. (b) Inference Phase) Routing: Given a test image, the system first identifies the most suitable category by minimizing the accumulated nearest-prototype distance across all candidate memory banks, referred to as Geometry-Consistent Routing. Scoring: After routing, the anomaly map is computed solely within the selected expert using patch-wise distances to the retrieved prototypes. This explicit separation between routing and scoring ensures stable anomaly prediction under task-agnostic, multi-category, and continual settings.
2 Related Work
2.1 Industrial Anomaly Detection
Industrial Anomaly Detection (IAD) aims to detect and localize visual defects under the constraint that only normal samples are available during training. This setting reflects real-world manufacturing scenarios and is typically addressed in unsupervised or weakly supervised regimes. Since the introduction of the MVTec AD benchmark, most studies have focused not only on image-level anomaly classification but also on pixel-level defect localization, where the representation of normal data and the design of anomaly scoring functions play a critical role.[bergmann2019mvtec] As the field evolves, the focus has expanded to more complex challenges through diverse datasets: VisA introduces scenarios with multiple instances and complex structures to overcome the limitations of well-aligned object views[zou2022spot], while MVTec LOCO-AD shifts the paradigm toward ”logical anomalies” where global context, such as component counts or spatial relationships, must be understood beyond simple local textures.[bergmann2022beyond] Furthermore, the MVTec 3D-AD dataset incorporates depth information to address defects identifiable only in three-dimensional space[bergmann2021mvtec], and the large-scale Real-IAD benchmark provides a more robust evaluation under varying illumination and viewing angles, pushing the boundaries of IAD toward more practical and scalable industrial applications[wang2024real].
Existing approaches can be broadly categorized by how normality is defined and modeled. Early works predominantly adopted reconstruction-based methods, where anomalies are identified through reconstruction errors produced by autoencoders or related generative models. Representative examples include autoencoder variants enhanced with structural similarity measures, self-supervised segmentation networks, and discriminatively trained reconstruction embeddings. While conceptually intuitive, these methods often struggle to preserve fine-grained textures and may generalize poorly to complex industrial patterns.[bergmann2018improving, song2021anoseg, zavrtanik2021draem]Moreover, anomaly scores in these approaches are typically defined through reconstruction errors or likelihood-related quantities, which are used directly without explicit normalization across samples or evaluation settings.
Subsequently, probabilistic density estimation approaches emerged, explicitly modeling the distribution of normal data in feature space.[rudolph2021same, gudovskiy2022cflow, yu2021fastflow, zhang2023unsupervised] Normalizing-flow-based methods estimate likelihoods for anomaly scoring and have demonstrated strong performance with pixel-level localization, benefiting from exact density estimation and invertible mappings. However, their training complexity and sensitivity to architectural choices remain non-trivial challenges. More recently, diffusion-based models have been explored for anomaly detection by reconstructing normal patterns or removing defects through iterative denoising processes.[wyatt2022anoddpm, zhang2025diffusionad] Extensions incorporating text conditions or prompt-driven generation further expand this line of research toward controllable anomaly synthesis and augmentation.[sun2025unseen]Although these methods provide principled probabilistic formulations, the resulting anomaly scores are often applied directly, with limited consideration of their comparability across different categories or heterogeneous feature manifolds.
In parallel, feature-based anomaly detection leveraging powerful pre-trained vision encoders has become the dominant paradigm in industrial inspection. Instead of reconstructing inputs, these methods distinguish normal and anomalous samples directly in feature space by modeling the distribution or structural relationships of normal features. This category includes teacher–student knowledge distillation frameworks[wang2021student, deng2022anomaly, gu2023remembering], one-class classification methods,[zhou2021vae, yi2020patch, liu2023simplenet] explicit feature distribution modeling approaches, and memory-bank-based nearest-neighbor methods.[roth2022towards, bae2023pni] Owing to their strong performance and relatively simple pipelines, such approaches have inspired a wide range of follow-up studies.However, anomaly scores in these methods are typically defined as unnormalized distances or similarity measures in feature space, implicitly assuming that such scores remain meaningful across different local neighborhoods or categories.
Recently, there has been growing interest in extending unsupervised feature-based anomaly detection to unified multi-class and continual learning settings, where multiple categories are handled jointly or introduced sequentially.[you2022unified, guo2025dinomaly] At the same time, vision–language models and large foundation models have been adopted to enable zero-shot or prompt-based anomaly detection, significantly broadening the scope of industrial inspection beyond fixed-category assumptions.[jeong2023winclip, zhou2023anomalyclip, li2024promptad, gu2024anomalygpt] These trends indicate a shift toward more flexible and scalable anomaly detection systems.While these works emphasize system-level scalability and robustness, most operate on top of predefined anomaly scoring functions, without explicitly revisiting how anomaly scores themselves are defined or normalized under heterogeneous conditions.
Despite this progress, many anomaly detection pipelines in heterogeneous settings still rely on implicit assumptions about how decisions are made across multiple heads or tasks. While substantial effort has been devoted to feature representations, memory designs, and learning strategies, comparatively less attention has been paid to the decision rule used when an input must be assigned to one of several category-specific normality models without task labels. In multi-category and continual anomaly detection, this issue becomes particularly salient because overall performance can depend strongly on the stability of cross-head routing decisions under heterogeneous normal manifolds.
A common practice in task-agnostic inference is to route by directly comparing head-specific anomaly scores, where each head is constructed independently and may induce different score scales and tail behaviors. Consequently, even when the underlying feature space separates categories reasonably well, small cross-head mismatches in score distributions can lead to inconsistent expert selection, which in turn can dominate downstream anomaly detection and localization outcomes. This provides a complementary explanation for the apparent performance collapse sometimes observed in task-agnostic continual evaluation, beyond representational considerations.
In this work, we revisit task-agnostic continual anomaly detection from the perspective of cross-head decision making. Rather than introducing additional representation learning objectives or task-specific continual learning mechanisms, we focus on designing a routing rule that operates in a shared frozen feature space and does not require direct cross-head comparability of head-specific anomaly scores. This perspective reframes a portion of the continual failure mode as a decision stability problem in routing, motivating geometry-consistent expert selection as a simple and effective design principle.
3 Method
3.1 Setup and Notation
We consider unsupervised anomaly detection for industrial inspection under a category-incremental (continual) setting. An input image is $x\in\mathbb{R}^{3\times H_{0}\times W_{0}}$. Training data contains only normal images, and categories arrive sequentially. Let $\mathcal{C}$ denote the set of all categories and $|\mathcal{C}|=C_{\text{cat}}$.
At test time, the category identity is unknown and we output (i) a pixel-level anomaly map $M(x)\in\mathbb{R}^{H_{0}\times W_{0}}$ and (ii) an image-level anomaly score $\mathrm{score}(x)\in\mathbb{R}$. We assume the set of candidate categories available at a given stage is the set of previously observed categories (details in the experimental protocol). This setting can be viewed as a multi-head inference problem: each observed category induces its own head, while the system must select a plausible head at inference time without labels. Accordingly, we separate (i) a routing rule for head selection and (ii) a within-head scoring rule for anomaly localization.
3.2 Frozen Multi-Layer Patch Features
We extract patch-level features using a frozen pre-trained vision encoder $f(\cdot)$. In our implementation, $f$ is the visual tower of OpenCLIP (ViT) and we record intermediate token embeddings via forward hooks. Let $\mathcal{L}$ be a set of transformer block indices at which we record token embeddings. For each $\ell\in\mathcal{L}$, we obtain
$Z^{(\ell)}\in\mathbb{R}^{B\times(1+N)\times C_{\ell}},$
where the first token corresponds to [CLS] and the remaining $N$ tokens correspond to patch embeddings. We discard [CLS] and reshape the $N$ patch tokens into a $H^{\prime}\times W^{\prime}$ grid ($N=H^{\prime}W^{\prime}$), yielding a feature map
$F^{(\ell)}\in\mathbb{R}^{B\times C_{\ell}\times H^{\prime}\times W^{\prime}}.$
We concatenate multi-layer features along the channel dimension:
$F=\mathrm{Concat}_{\ell\in\mathcal{L}}F^{(\ell)}\in\mathbb{R}^{B\times D\times H^{\prime}\times W^{\prime}},\qquad D=\sum_{\ell\in\mathcal{L}}C_{\ell}.$ (1)
For a single image $x$ (dropping the batch index), we denote the patch feature at location $(u,v)$ by $q_{u,v}(x)\in\mathbb{R}^{D}$, and also use a flattened index $p\in\{1,\dots,N\}$ to write $q_{p}(x)\in\mathbb{R}^{D}$. The encoder $f$ is kept frozen throughout, and we operate directly on these patch embeddings. Multi-layer concatenation is used to combine intermediate representations at different levels of abstraction without introducing additional learned transformations.
3.3 Per-Category Prototype Bank via Coreset Selection
For each category $c\in\mathcal{C}$, we construct a prototype bank
$\mathcal{M}_{c}=\{\mu_{c,1},\dots,\mu_{c,K}\},\qquad\mu_{c,k}\in\mathbb{R}^{D}.$ (2)
Let $\mathcal{Q}_{c}=\{q_{p}(x)\,|\,x\in\mathcal{D}^{\mathrm{train}}_{c},\,p=1,\dots,N\}$ denote the multiset of all patch features from normal training images of category $c$. We select $K$ representatives from $\mathcal{Q}_{c}$ using greedy $k$-center (farthest-first) coreset selection under squared Euclidean distance in feature space:
$\mu_{c,1}\sim\mathrm{Uniform}(\mathcal{Q}_{c}),\qquad\mu_{c,t}=\arg\max_{q\in\mathcal{Q}_{c}}\ \min_{1\leq s<t}\ \|q-\mu_{c,s}\|_{2}^{2},\quad t=2,\dots,K.$ (3)
Intuitively, $\mathcal{M}_{c}$ provides a compact geometric summary of normal patch embeddings for category $c$. The farthest-first rule encourages coverage of $\mathcal{Q}_{c}$ under the chosen metric, which is beneficial both for within-head scoring and for routing across heads. We implement Eq. (3) by maintaining the current nearest-prototype distance for each candidate feature and updating it after each new selection. The resulting bank $\mathcal{M}_{c}$ is stored in the original $D$-dimensional space.
3.4 Prototype-Based Patch Scoring (Within-Head)
Given a patch feature $q\in\mathbb{R}^{D}$ and a category head $c$ with prototypes $\{\mu_{c,k}\}_{k=1}^{K}$, we compute a patch-level anomaly score using distances to prototypes. We first define squared distances
$d_{c,k}(q)=\|q-\mu_{c,k}\|_{2}^{2}.$ (4)
This distance-based scoring is intentionally lightweight: it does not introduce additional learnable parameters and can be instantiated with either hard nearest-prototype matching or a smooth soft-min aggregation.
We aggregate distances across prototypes with a soft-min (LogSumExp) operator:
$E_{c}(q)=-\log\sum_{k=1}^{K}\pi_{c,k}\,\exp\!\left(-\frac{1}{2\tau}\,d_{c,k}(q)\right),$ (5)
where $\pi_{c,k}\geq 0$, $\sum_{k}\pi_{c,k}=1$ (uniform by default), and $\tau>0$ is a temperature controlling the softness of aggregation. As $\tau\to 0$, Eq. (5) approaches the hard-min form (up to constants), concentrating on the nearest prototype. Optionally, we restrict the summation in Eq. (5) to the top-$K_{e}$ nearest prototypes to reduce computation.
For an image $x$ with patch grid $H^{\prime}\times W^{\prime}$, we define the category-specific patch anomaly map $S_{c}(x)\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ as
$S_{c}(x)[u,v]:=E_{c}\bigl(q_{u,v}(x)\bigr).$ (6)
We use $S_{c}(x)$ only after a head has been selected; within-head scoring is not required to be comparable across heads for routing.
3.5 Task-Agnostic Routing via Geometry-Consistent Prototype Matching
At inference time, the category identity is unknown. Given candidate categories $\mathcal{C}_{\mathrm{cand}}\subseteq\mathcal{C}$ available at the current stage, we perform routing directly in the frozen patch-embedding space using per-category prototype banks. The routing objective is to select the head whose prototype bank best matches the overall patch geometry of the input under a shared metric, without relying on direct comparability of head-specific anomaly scores.
We define a routing distance for each $c\in\mathcal{C}_{\mathrm{cand}}$ by accumulating patch-wise nearest-prototype squared distances:
$r_{c}(x)=\sum_{p=1}^{N}\min_{k\in\{1,\dots,K\}}\bigl\|q_{p}(x)-\mu_{c,k}\bigr\|_{2}^{2}.$ (7)
Equation (7) aggregates nearest-prototype distances over patches, so routing depends on global consistency of patch-to-prototype matches rather than on a small number of highly deviant patches. In practice, we approximate Eq. (7) by evaluating a random subset of $M\ll N$ patches for efficiency; this does not change the definition of the routing rule.
We route the input to the closest category (top-1 gating):
$\hat{c}(x)=\arg\min_{c\in\mathcal{C}_{\mathrm{cand}}}r_{c}(x).$ (8)
3.6 Final Anomaly Map and Image Score
Given the routed category $\hat{c}(x)$, we compute a single anomaly map
$S(x):=S_{\hat{c}(x)}(x)\in\mathbb{R}^{H^{\prime}\times W^{\prime}}.$ (9)
We upsample $S(x)$ to the input resolution $(H_{0},W_{0})$ using bilinear interpolation to obtain $M(x)\in\mathbb{R}^{H_{0}\times W_{0}}$. For the image-level anomaly score, we aggregate the upsampled map by a top-$q$ pooling operator:
$\mathrm{score}(x):=\mathrm{TopQPool}\bigl(M(x)\bigr),$ (10)
where $\mathrm{TopQPool}$ averages the largest $q$ fraction of pixel scores (we use $q=1\%$ by default). Top-$q$ pooling emphasizes concentrated defect regions and is less sensitive to diffuse background responses.
4 Experiment
4.1 Experimental Setup
Datasets.
We evaluate GCR on MVTec AD, which contains 15 categories (objects and textures). Following the standard protocol, only normal images are used for training, while the test set contains both normal and anomalous samples. We report both image-level anomaly detection and pixel-level anomaly localization.
Feature extractor.
We use a frozen ViT-B/16 from OpenCLIP/CLIP [schuhmann2021laion]. For each input image, we extract multi-layer patch token features by recording intermediate outputs at a fixed set of transformer blocks and concatenating them along the channel dimension. The encoder is kept frozen in all experiments; all routing and anomaly scoring are performed directly in the resulting patch embedding space.
Baselines and comparison protocol.
We compare GCR with representative anomaly detection methods, including probabilistic modeling approaches (PaDiM [defard2021padim]), memory-based retrieval methods (PatchCore [roth2022towards], CFA [lee2022cfa]), and recent unified or continual anomaly detection frameworks (SimpleNet [liu2023simplenet], UniAD [you2022unified], UCAD [liu2024unsupervised]). For methods originally designed for category-aware inference, we follow their official implementations and evaluation protocols. Unless otherwise stated, all baselines are evaluated under the same backbone, input resolution, and dataset splits. For continual experiments, we follow the standard sequential protocol and report forgetting using image-level AUROC.
Prototype bank construction (per-category coreset).
For each category $c$, we construct a prototype bank $\mathcal{M}_{c}=\{\mu_{c,k}\}_{k=1}^{K}$ from normal training images. We first collect all patch features from the training set of category $c$, and then select $K$ representative prototypes using greedy $k$-center (farthest-first) coreset selection under squared Euclidean distance. Unless otherwise stated, prototypes are fixed after selection and are not updated by learning.
Scoring and inference pipeline (task-agnostic).
At test time, the category identity is assumed unknown. We perform task-agnostic routing directly in the patch embedding space. For each category $c$, we compute a routing distance by accumulating patch-wise nearest-prototype distances:
$r_{c}(x)=\frac{1}{N}\sum_{p=1}^{N}\min_{k}\|q_{p}(x)-\mu_{c,k}\|_{2}^{2}.$
We route each test sample to the top-1 category $\hat{c}(x)=\arg\min_{c}r_{c}(x)$ by default. Given the routed head, we compute a patch anomaly map $S_{\hat{c}}(x)$ using our prototype-based scoring rule (LSE energy by default, with hard-min as a comparable variant). The anomaly map is upsampled to the input resolution by bilinear interpolation, and the image-level score is computed by top-$q$ pooling over pixels. Optionally, one may fuse the top-$k$ routed heads by an elementwise maximum; we treat this as a robustness option rather than the default.
Hyperparameters.
We use a single set of hyperparameters across all categories. The default settings in our main experiments are: topk=1 for routing and topq=0.01 for image-level pooling. For patch scoring, we use a smooth LogSumExp (LSE) aggregation over prototypes by default; hard-min nearest-prototype scoring yields comparable rank-based metrics in our setting and is included in ablations. Unless explicitly varied, we use uniform component weights $\pi_{c,k}=1/K$ and keep all computations in the frozen patch embedding space.
Implementation details.
All experiments are run on a single NVIDIA RTX 4090 GPU. Prototype selection is performed once per category offline. At inference, the dominant computation is the encoder forward pass and the evaluation of prototype energies for the routed candidate heads.
4.2 Evaluation Metrics
For each test image $x$, the method outputs (i) an image-level anomaly score $\mathrm{score}(x)\in\mathbb{R}$ and (ii) a pixel-level anomaly map $M(x)\in\mathbb{R}^{H_{0}\times W_{0}}$. The anomaly map is upsampled to the input resolution by bilinear interpolation. Unless stated otherwise, the image-level score is obtained by top-$q$ pooling over pixels (i.e., the mean of the largest $q$ fraction of values in $M(x)$), which is robust to localized defects.
Image-level detection.
We report image-level AUROC and Average Precision (AP). Given test pairs $\{(\mathrm{score}(x_{i}),y_{i})\}_{i=1}^{n}$ with $y_{i}\in\{0,1\}$, AUROC is threshold-free and measures ranking quality between normal and anomalous images. AP is the area under the precision–recall (PR) curve and is particularly informative under class imbalance.
Pixel-level localization.
For pixel-level localization, each pixel is treated as a binary classification instance using the anomaly map $M(x)$ and the ground-truth mask $m(x)\in\{0,1\}^{H_{0}\times W_{0}}$. We report pixel-level AUROC (p-AUROC) and pixel-level AP, denoted as p-AP (equivalently p-AUPRC/p-AUPC). Let $\{(s_{\ell},t_{\ell})\}_{\ell=1}^{L}$ be the set of pixel scores and labels aggregated over the test set, where $s_{\ell}$ is a pixel score and $t_{\ell}\in\{0,1\}$ indicates anomaly. p-AUROC evaluates separability of anomalous vs. normal pixels across thresholds. p-AP summarizes the PR curve and penalizes false positives more directly, which is crucial since anomalous regions typically occupy only a small fraction of pixels.
Routing diagnostics (task-agnostic inference).
In task-agnostic evaluation, overall detection performance depends on whether a test sample is routed to an appropriate category head. We therefore report routing accuracy
$\mathrm{Acc}_{\mathrm{route}}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\bigl[\hat{c}(x_{i})=c_{i}\bigr],$
where $c_{i}$ is the ground-truth category and $\hat{c}(x_{i})$ is the routed category. To isolate the effect of routing errors, we additionally report conditional image-level AUROC on the subsets $\{\hat{c}=c\}$ and $\{\hat{c}\neq c\}$. These diagnostics help separate routing failures from within-head scoring failures.
Continual evaluation and forgetting measure.
We adopt the standard forgetting measure (FM) in continual learning [chaudhry2018riemannian]. We consider a sequential setting where categories arrive in order $t=1,\dots,T$. At step $t$, we build a new head (prototype bank and associated scoring parameters) for the new category and keep previously constructed heads unchanged. After each step $t$, we evaluate the system on each previously seen category $i\leq t$ using the task-agnostic inference pipeline (i.e., routing among the currently available heads; optional top-$k$ fusion is reported separately when used), and compute a performance score $P_{i}^{(t)}$. The forgetting for category $i$ is
$\mathrm{FM}_{i}=\max_{t\in\{i,\dots,T-1\}}P_{i}^{(t)}-P_{i}^{(T)},$
and the overall forgetting is
$\mathrm{FM}=\frac{1}{T-1}\sum_{i=1}^{T-1}\mathrm{FM}_{i}.$
In this work we use image-level AUROC as $P_{i}^{(t)}$ when computing FM. We emphasize that GCR does not update a shared representation to prevent forgetting. Instead, heads remain category-specific, and FM primarily reflect
Table 1: Continual multi-category anomaly detection on MVTec AD (Image-level AUROC). Higher is better. AvgFM is lower is better.
Class CFA [lee2022cfa] PaDiM [defard2021padim] PatchCore [roth2022towards] SimpleNet [liu2023simplenet] UniAD [you2022unified] UCAD [liu2024unsupervised] GCR (ours)
Bottle 0.309 0.458 0.533 0.938 0.997 1.000 1.000
Cable 0.489 0.544 0.505 0.560 0.701 0.751 0.975
Capsule 0.275 0.418 0.351 0.519 0.765 0.866 0.957
Carpet 0.834 0.454 0.865 0.736 0.998 0.965 1.000
Grid 0.571 0.704 0.723 0.592 0.896 0.944 1.000
Hazelnut 0.903 0.635 0.959 0.859 0.936 0.994 1.000
Leather 0.935 0.418 0.854 0.749 1.000 1.000 1.000
Metal Nut 0.464 0.446 0.456 0.710 0.964 0.988 0.998
Pill 0.528 0.449 0.511 0.701 0.895 0.894 0.958
Screw 0.528 0.578 0.626 0.599 0.554 0.739 0.855
Tile 0.763 0.581 0.748 0.654 0.989 0.998 1.000
Toothbrush 0.519 0.678 0.600 0.422 0.928 1.000 0.975
Transistor 0.320 0.407 0.427 0.669 0.966 0.874 0.977
Wood 0.923 0.549 0.900 0.908 0.982 0.995 0.995
Zipper 0.984 0.855 0.974 0.996 0.987 0.938 0.986
Average $\uparrow$ 0.623 0.545 0.669 0.708 0.904 0.930 0.978
AvgFM $\downarrow$ 0.361 0.368 0.318 0.211 0.076 0.010 0.000
Table 2: Pixel-level localization on MVTec AD (p-AP). Higher is better. AvgFM is lower is better. Bold indicates the best method per row.
Class CFA [lee2022cfa] PaDiM [defard2021padim] PatchCore [roth2022towards] SimpleNet [liu2023simplenet] UniAD [you2022unified] UCAD [liu2024unsupervised] GCR (ours)
Bottle 0.068 0.072 0.087 0.108 0.734 0.752 0.700
Cable 0.056 0.037 0.043 0.045 0.232 0.290 0.552
Capsule 0.050 0.030 0.042 0.029 0.313 0.349 0.336
Carpet 0.271 0.023 0.407 0.018 0.517 0.622 0.733
Grid 0.004 0.006 0.003 0.004 0.204 0.187 0.334
Hazelnut 0.341 0.183 0.443 0.029 0.378 0.506 0.530
Leather 0.393 0.039 0.352 0.006 0.360 0.333 0.423
Metal Nut 0.255 0.155 0.189 0.227 0.587 0.775 0.668
Pill 0.080 0.044 0.058 0.077 0.346 0.634 0.582
Screw 0.015 0.014 0.017 0.004 0.035 0.214 0.226
Tile 0.155 0.065 0.124 0.082 0.428 0.549 0.598
Toothbrush 0.053 0.044 0.028 0.046 0.398 0.298 0.437
Transistor 0.056 0.049 0.053 0.049 0.542 0.398 0.452
Wood 0.281 0.080 0.270 0.037 0.378 0.535 0.552
Zipper 0.573 0.452 0.604 0.139 0.443 0.398 0.575
Average $\uparrow$ 0.177 0.086 0.181 0.060 0.393 0.456 0.513
AvgFM $\downarrow$ 0.083 0.366 0.343 0.069 0.086 0.013 0.000
Table 3: Continual multi-category anomaly detection on VisA (Image-level AUROC). Higher is better. AvgFM is lower is better.
Class CFA [lee2022cfa] PatchCore[roth2022towards] SimpleNet [liu2023simplenet] UniAD [you2022unified] UCAD [liu2024unsupervised] GCR (ours)
Candle 0.512 0.647 0.504 0.573 0.778 0.910
Capsules 0.672 0.579 0.474 0.599 0.877 0.788
Cashew 0.873 0.669 0.794 0.661 0.960 0.957
Chewinggum 0.753 0.735 0.721 0.758 0.958 0.972
Fryum 0.304 0.431 0.684 0.504 0.945 0.926
Macaroni1 0.557 0.631 0.567 0.559 0.823 0.824
Macaroni2 0.422 0.624 0.447 0.644 0.667 0.766
PCB1 0.698 0.617 0.598 0.749 0.905 0.942
PCB2 0.472 0.534 0.629 0.523 0.871 0.844
PCB3 0.449 0.479 0.538 0.547 0.813 0.808
PCB4 0.407 0.645 0.493 0.562 0.901 0.976
PipeFryum 0.998 0.999 0.945 0.989 0.988 1.000
Average $\uparrow$ 0.593 0.633 0.616 0.639 0.874 0.893
AvgFM $\downarrow$ 0.327 0.349 0.283 0.297 0.039 0.000
Table 4: Continual multi-category anomaly localization on VisA (pixel-level p-AP / p-AUPRC / p-AUPC). Higher is better. AvgFM is lower is better. Bold indicates the best method per row.
Class CFA [lee2022cfa] PatchCore [roth2022towards] SimpleNet [liu2023simplenet] UniAD [you2022unified] UCAD [liu2024unsupervised] GCR (ours)
Candle 0.017 0.018 0.001 0.132 0.067 0.135
Capsules 0.005 0.010 0.004 0.123 0.437 0.386
Cashew 0.059 0.047 0.017 0.378 0.580 0.457
Chewinggum 0.243 0.202 0.007 0.574 0.503 0.558
Fryum 0.085 0.081 0.047 0.404 0.334 0.281
Macaroni1 0.001 0.003 0.000 0.041 0.013 0.069
Macaroni2 0.001 0.001 0.000 0.010 0.003 0.041
PCB1 0.013 0.008 0.013 0.612 0.702 0.652
PCB2 0.006 0.004 0.003 0.083 0.136 0.107
PCB3 0.008 0.008 0.004 0.266 0.266 0.253
PCB4 0.015 0.010 0.009 0.232 0.106 0.319
PipeFryum 0.592 0.443 0.058 0.549 0.457 0.367
Average $\uparrow$ 0.087 0.070 0.014 0.283 0.300 0.302
AvgFM $\downarrow$ 0.184 0.327 0.016 0.062 0.015 0.000
Table 1 summarizes continual multi-category results on MVTec AD under task-agnostic inference. GCR achieves the highest average image-level AUROC and competitive pixel-level localization among compared methods. We also observe near-zero forgetting in our evaluation, suggesting that performance on previously seen categories remains stable as new categories are introduced.
Notably, these gains are obtained without end-to-end representation learning. Because task identity is unknown at test time, incorrect routing can dominate performance even when within-head scoring is strong. In the following ablations, we disentangle (i) the routing decision rule and (ii) the within-head scoring form, and show that geometry-consistent routing is the primary contributor to the observed improvements.
Figure 2: Qualitative anomaly localization results of GCR on MVTec AD. For each example, we show the input image (left) and the corresponding anomaly map (right), obtained by task-agnostic routing, optional top-$k$ head fusion, and top-$q$ pooling. GCR consistently highlights defect regions while suppressing irrelevant background responses across both object and texture categories, including subtle and spatially localized anomalies.
4.3 Ablation Study
We present ablations under the continual multi-category setting with task-agnostic inference. The goal of this section is to identify which factor primarily determines continual performance: (i) the routing decision rule across category heads, or (ii) the within-head anomaly scoring form and its optional variants. This distinction is crucial because in task-agnostic continual evaluation, a test sample must first be assigned to a plausible head before any head-specific scoring can be meaningful.
Routing diagnostics and conditional evaluation.
In addition to overall AUROC, we report routing accuracy $\mathrm{Acc}_{\mathrm{route}}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}[\hat{c}(x_{i})=c_{i}]$, where $c_{i}$ is the ground-truth category and $\hat{c}(x_{i})$ is the routed category. To isolate the impact of routing errors, we also report conditional image-level AUROC on the subsets $\{\hat{c}=c\}$ and $\{\hat{c}\neq c\}$.
Ablation I: Routing rule is the dominant factor.
We first ablate the routing decision rule while keeping the backbone frozen and using identical per-category prototype banks. We compare: (a) score-based routing, which selects the head by comparing head-specific anomaly scores (e.g., $\arg\min_{c}\,\mathrm{score}_{c}(x)$), and (b) geometry-consistent routing, which selects the head by accumulated nearest-prototype distances in the patch embedding space,
$\hat{c}(x)=\arg\min_{c\in\mathcal{C}_{t}}r_{c}(x),\qquad r_{c}(x)=\frac{1}{N}\sum_{p=1}^{N}\min_{k}\|q_{p}(x)-\mu_{c,k}\|_{2}^{2},$
where $\mathcal{C}_{t}$ is the set of available heads at continual step $t$ and $N$ is the number of patch tokens.
Ablation I: Routing rule is the dominant factor.
We first ablate the routing decision rule while keeping the backbone frozen and using identical per-category prototype banks. We compare: (a) score-based routing, which selects the head by comparing head-specific anomaly scores (e.g., $\arg\min_{c}\,\mathrm{score}_{c}(x)$), and (b) geometry-consistent routing, which selects the head by accumulated nearest-prototype distances in the normalized patch embedding space,
$\hat{c}(x)=\arg\min_{c\in\mathcal{C}_{t}}r_{c}(x),\qquad r_{c}(x)=\frac{1}{N}\sum_{p=1}^{N}\min_{k}\|q_{p}(x)-\mu_{c,k}\|_{2}^{2},$
where $\mathcal{C}_{t}$ is the set of available heads at continual step $t$ and $N$ is the number of patch tokens. Table 5 reports routing accuracy (overall / normal / anomalous) together with image-level AUROC and pixel-level p-AP. We observe that geometry-consistent routing achieves perfect routing accuracy for both normal and anomalous samples, while score-based routing makes non-trivial routing errors particularly on anomalous inputs. This supports the view that continual performance degradation under task-agnostic inference is primarily a routing mismatch issue rather than a lack of representational capacity.
Table 5: Routing rule ablation under continual task-agnostic evaluation (MVTec AD, final step). We report routing accuracy (overall / normal / anomalous), image-level AUROC, and pixel-level p-AP.
Metric Score-based ($\arg\min_{c}\,\mathrm{score}_{c}(x)$) Geometry-consistent ($\arg\min_{c}\,r_{c}(x)$)
Routing Acc. $\uparrow$ 0.937 1.000
Routing Acc. (Normal) $\uparrow$ 1.000 1.000
Routing Acc. (Anomaly) $\uparrow$ 0.914 1.000
AUROC (%) $\uparrow$ 98.08 98.08
p-AP $\uparrow$ 0.380 0.508
Ablation II: Scoring-form invariance given fixed routing.
Next, we fix routing to the geometry-consistent rule and study the effect of within-head scoring parameterizations. We compare energy-based scoring and Gaussian negative log-likelihood (NLL), as well as hard vs. soft aggregation over prototypes.
Let $q\in\mathbb{R}^{D}$ denote a patch feature and $\{\mu_{c,k}\}_{k=1}^{K}$ denote prototypes for head $c$. Under diagonal precisions $\Lambda_{c,k}=\mathrm{diag}(\lambda_{c,k})$ and mixture weights $\pi_{c,k}$, energy-based scoring can be written as a LogSumExp aggregation of (negative) quadratic forms:
$E_{c}(q)=-\log\sum_{k=1}^{K}\pi_{c,k}\,\exp\!\left(-\frac{1}{2}\Bigl[(q-\mu_{c,k})^{\top}\Lambda_{c,k}(q-\mu_{c,k})-\log|\Lambda_{c,k}|\Bigr]\right).$
A Gaussian NLL differs from this expression by additive constants and, under fixed/shared precisions, yields an equivalent ordering over patches. Similarly, hard-min nearest-prototype scoring and soft LogSumExp aggregation are closely related; as temperature decreases, LSE approaches the hard minimum. Since AUROC depends only on score ranking, such equivalent forms are expected to produce similar results.
Table 6 confirms that, once routing is fixed, energy and NLL yield nearly identical AUROC and FM, and min vs. LSE aggregation produces comparable results. This indicates that our gains do not hinge on a particular probabilistic parameterization or aggregation operator, but rather on the task-agnostic routing rule that determines which head is evaluated.
Table 6: Scoring-form invariance given geometry-consistent routing in continual evaluation (MVTec AD). Image-level AUROC (%) and forgetting measure (FM) are reported. Energy vs. NLL and Min vs. LSE yield similar rank-based performance; adaptation is optional and can reduce stability.
Scoring Aggregation Adaptation AUROC / FM
Energy LSE Yes 97.7 / 0.001
Energy LSE No 97.9 / 0.000
Energy Min Yes 97.8 / 0.008
Energy Min No 97.5 / 0.000
NLL LSE Yes 97.7 / 0.001
NLL LSE No 97.9 / 0.000
NLL Min Yes 97.8 / 0.008
NLL Min No 97.5 / 0.000
Ablation III: Effect of lightweight adaptation.
We evaluate lightweight adaptation of geometric parameters while keeping the backbone encoder frozen. In our implementation, adaptation updates a small set of diagonal log-precision parameters (per head) using an exponential moving average (EMA) estimate of per-prototype variance on normal training patches. While such adaptation can be interpreted as fitting head-specific local uncertainty, it also introduces additional degrees of freedom that may increase cross-head heterogeneity under task-agnostic routing.
On MVTec AD, Table 6 shows that enabling adaptation does not improve AUROC and can increase FM. the training-free configuration achieves both higher routing accuracy and markedly improved macro AUROC, while eliminating forgetting. These results support our design choice to treat adaptation as an auxiliary option rather than a core requirement.
Ablation IV: Backbone feature layer selection.
Finally, we study the effect of the selected transformer block for patch feature extraction. Table 7 reports macro image-level AUROC over 15 categories on MVTec AD. Mid-level layers (5–6) yield the strongest macro performance in our setup, and we use layer index 6 by default. We note that while feature layer choice affects absolute performance, it does not change the main conclusion of this section: routing stability remains the dominant driver of task-agnostic continual performance once a reasonable frozen feature representation is used.
Table 7: Effect of backbone feature layer selection. Macro image-level AUROC is reported over 15 categories on MVTec AD.
Layer index AUROCmacro (%)
4 96.33
5 97.87
6 97.93
7 97.48
8 97.11
9 96.87
5 Conclusion
We revisited task-agnostic continual anomaly detection from a decision-rule perspective and argued that a major source of performance collapse is not necessarily representational insufficiency, but the head-selection rule used when task identity is unknown. In multi-head continual settings, routing is inherently a cross-head comparison problem: the system must decide which category head to trust before applying any head-specific anomaly model. Our experiments show that naïvely routing by comparing heterogeneous anomaly scores across independently constructed heads is ill-posed and can dominate the overall failure mode.
Based on this observation, we proposed GCR, a geometric normalization view of anomaly scoring that emphasizes geometry-consistent routing in a shared patch-embedding space. Concretely, we route test samples by accumulated nearest-prototype distances under $\ell_{2}$-normalized patch features, and then perform within-head anomaly localization using prototype-based scoring. Crucially, once routing is stabilized in a shared geometric criterion, the exact probabilistic parameterization of within-head scoring (energy vs. NLL, hard-min vs. LSE aggregation) has only a minor effect on ranking-based metrics, while optional lightweight adaptation can reduce stability under continual evaluation.
Across MVTec AD and VisA, GCR achieves strong task-agnostic continual performance with near-zero forgetting, without end-to-end representation learning. More broadly, our results suggest that continual failures previously attributed to catastrophic forgetting can, in part, be reframed as decision instability caused by mismatched cross-head score scales. This points to a complementary research direction: improving continual anomaly detection by designing routing and scoring rules that remain comparable across heterogeneous normal manifolds, rather than relying solely on additional feature separation objectives.
Limitations and future work.
Our current framework assumes that each category head maintains its own prototype bank, which scales linearly with the number of categories. Future work should investigate memory-sharing strategies while preserving geometry-consistent routing. In addition, while we focused on ranking-based metrics, a deeper study of score calibration across heads and operating-point stability (e.g., threshold transfer across stages) would further strengthen deployment-oriented continual AD systems.
Appendix A Notation and Dimensional Consistency
For clarity, we summarize the notation and dimensional conventions used throughout the paper, and explicitly distinguish between feature dimensions and category indices.
Input and categories.
An input image is denoted by $x\in\mathbb{R}^{3\times H_{0}\times W_{0}}$. The set of categories is $\mathcal{C}$ with cardinality $|\mathcal{C}|=C_{\mathrm{cat}}$. Category indices are denoted by $c\in\mathcal{C}$.
Patch features.
A frozen vision transformer encoder produces multi-layer patch features, which are concatenated to form a feature tensor
$F\in\mathbb{R}^{D\times H^{\prime}\times W^{\prime}},$
where $D$ denotes the feature dimension after concatenation, and $H^{\prime}\times W^{\prime}$ is the spatial resolution of patch tokens. We flatten spatial indices and denote the $p$-th patch feature by
$q_{p}(x)\in\mathbb{R}^{D},\qquad p=1,\dots,N,\quad N=H^{\prime}W^{\prime}.$
Prototype banks.
For each category $c$, a prototype bank
$\mathcal{M}_{c}=\{\mu_{c,1},\dots,\mu_{c,K}\},\qquad\mu_{c,k}\in\mathbb{R}^{D},$
is constructed from normal training images using greedy $k$-center coreset selection. The integer $K$ controls the geometric coverage of the normal feature manifold.
Geometric parameters.
Each prototype can optionally be associated with a diagonal precision vector
$\lambda_{c,k}\in\mathbb{R}^{D}_{+},$
which defines an anisotropic quadratic form in feature space. Unless explicitly enabled, we use $\lambda_{c,k}=\mathbf{1}$, corresponding to an isotropic metric.
This notation eliminates potential ambiguity between feature dimensions ($D$), category indices ($c$), and the number of categories ($C_{\mathrm{cat}}$).
Appendix B Probabilistic Interpretation and Limiting Cases
B.1 Energy as a Temperature-Scaled Gaussian Mixture Model
We show that the proposed energy admits a probabilistic interpretation as a temperature-scaled negative log-likelihood of a Gaussian mixture model (GMM) defined in feature space.
For a fixed category $c$, consider the mixture distribution
$p(q\mid c)=\sum_{k=1}^{K}\pi_{c,k}\,\mathcal{N}\!\left(q;\,\mu_{c,k},\,\Lambda_{c,k}^{-1}\right),$ (11)
where $\pi_{c,k}\geq 0$, $\sum_{k}\pi_{c,k}=1$, and $\Lambda_{c,k}=\mathrm{diag}(\lambda_{c,k})$ is a diagonal precision matrix.
The negative log-likelihood of $q$ under this model is
$\log p(q\mid c)=-\log\sum_{k=1}^{K}\pi_{c,k}\,\exp\!\left(-\frac{1}{2}\Bigl[(q-\mu_{c,k})^{\top}\Lambda_{c,k}(q-\mu_{c,k})-\log|\Lambda_{c,k}|\Bigr]\right)+\mathrm{const},$ (12)
where the constant term is independent of $q$.
Introducing a temperature parameter $T>0$ yields the scaled energy
$E_{c,T}(q)=-T\log\sum_{k=1}^{K}\pi_{c,k}\,\exp\!\left(-\frac{1}{2T}\Bigl[(q-\mu_{c,k})^{\top}\Lambda_{c,k}(q-\mu_{c,k})-\log|\Lambda_{c,k}|\Bigr]\right),$ (13)
which matches the main-text energy up to an additive constant independent of $q$. Therefore, the proposed score can be interpreted as a temperature-scaled negative log-likelihood under a locally defined GMM in feature space.
B.2 Relation to Location-Wise Gaussian Modeling
Location-wise Gaussian models such as PaDiM can be viewed as a restricted instance of this formulation.
In PaDiM, for each spatial location $(u,v)$ and category $c$, a single Gaussian distribution
$\mathcal{N}(q;\,\mu_{c,(u,v)},\,\Sigma_{c,(u,v)})$
is estimated from training data, and anomaly scores are computed using the corresponding Mahalanobis distance.
This corresponds to a degenerate case where: (i) the mixture index $k$ is replaced by a fixed spatial index, (ii) only one component is used per location, and (iii) component assignment is fixed rather than query-conditioned.
In contrast, our formulation uses multiple local components independently of spatial position and performs query-conditioned soft (or hard) assignment in feature space, which enables score computation relative to local feature geometry rather than fixed spatial statistics.
Appendix C Effect of Prototype Coverage
We analyze the effect of prototype coverage $K$ using a full K-sweep under task-agnostic inference. Unless stated otherwise, all experiments in this appendix satisfy: (i) the feature extractor is frozen, (ii) prototype banks are fixed after construction, (iii) no EMA update is applied unless explicitly enabled, and (iv) routing is performed without access to ground-truth category labels.
For each category, we report image-level AUROC under oracle inference (ground-truth head), routed inference (task-agnostic routing), and the corresponding forgetting measure (FM) in the continual setting.
C.1 EMA OFF: Training-Free Geometry Baseline
Tables 810 report the full K-sweep results when no EMA-based anisotropy estimation is applied (EMA OFF). In this setting, anomaly scores are computed purely from fixed prototype geometry.
As $K$ increases, we observe consistent improvements in both oracle and routed AUROC. Moreover, once moderate coverage is reached (typically $K\geq 64$ in our experiments), routed performance becomes close to oracle performance and FM rapidly approaches zero. This suggests that, in the considered setting, routing ambiguity and performance instability are largely mitigated by sufficient geometric coverage, without requiring additional parameter adaptation.
Table 8: EMA OFF. Oracle image-level AUROC across prototype coverage $K$. The last row reports the average over categories.
Category 16 32 64 96 128 196 256 512
Bottle .948 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Cable .730 .852 .924 .936 .946 .976 .979 .987
Capsule .736 .854 .896 .896 .921 .952 .968 .987
Carpet .909 .986 .997 1.00 1.00 1.00 1.00 1.00
Grid .779 .996 .990 .993 1.00 1.00 1.00 1.00
Hazelnut .956 .948 1.00 1.00 1.00 1.00 1.00 1.00
Leather 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
MetalNut .925 .983 .981 .990 .994 .998 1.00 1.00
Pill .853 .885 .949 .933 .941 .956 .956 .968
Screw .519 .495 .627 .719 .720 .856 .886 .947
Tile .954 1.00 1.00 .999 1.00 1.00 1.00 1.00
Toothbrush .783 .997 .969 .994 .983 .972 .967 .983
Transistor .584 .855 .959 .993 .975 .977 .973 .989
Wood .956 1.00 .996 .995 .993 .995 .994 .996
Zipper .879 .962 .972 .970 .984 .987 .988 .992
Mean .834 .921 .951 .961 .964 .978 .981 .990
Table 9: EMA OFF. Routed image-level AUROC across prototype coverage $K$. The last row reports the average over categories.
Category 16 32 64 96 128 196 256 512
Bottle .948 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Cable .730 .852 .924 .936 .946 .976 .979 .987
Capsule .736 .854 .896 .896 .921 .952 .968 .987
Carpet .909 .986 .997 1.00 1.00 1.00 1.00 1.00
Grid .779 .996 .990 .993 1.00 1.00 1.00 1.00
Hazelnut .828 .734 1.00 1.00 1.00 1.00 1.00 1.00
Leather 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
MetalNut .925 .983 .981 .990 .994 .998 1.00 1.00
Pill .590 .885 .949 .933 .941 .956 .956 .968
Screw .519 .495 .627 .719 .720 .856 .886 .947
Tile .954 1.00 1.00 .999 1.00 1.00 1.00 1.00
Toothbrush .789 .997 .969 .994 .983 .972 .967 .983
Transistor .572 .855 .959 .993 .975 .977 .973 .989
Wood .956 1.00 .996 .995 .993 .995 .994 .996
Zipper .879 .962 .972 .970 .984 .987 .988 .992
Mean .808 .907 .951 .961 .964 .978 .981 .990
Table 10: EMA OFF. Forgetting measure (FM) across prototype coverage $K$. The last row reports the average over categories.
Category 16 32 64 96 128 196 256 512
Bottle .001 .000 .000 .000 .000 .000 .000 .000
Cable .000 .000 .000 .000 .000 .000 .000 .000
Capsule .000 .000 .000 .000 .000 .000 .000 .000
Carpet .000 .000 .000 .000 .000 .000 .000 .000
Grid .000 .000 .000 .000 .000 .000 .000 .000
Hazelnut .009 .181 .000 .000 .000 .000 .000 .000
Leather .000 .000 .000 .000 .000 .000 .000 .000
MetalNut .000 .000 .000 .000 .000 .000 .000 .000
Pill .274 .000 .000 .000 .000 .000 .000 .000
Screw .000 .000 .000 .000 .000 .000 .000 .000
Tile .000 .000 .000 .000 .000 .000 .000 .000
Toothbrush .000 .000 .000 .000 .000 .000 .000 .000
Transistor .012 .000 .000 .000 .000 .000 .000 .000
Wood .000 .000 .000 .000 .000 .000 .000 .000
Zipper .000 .000 .000 .000 .000 .000 .000 .000
Mean .020 .012 .000 .000 .000 .000 .000 .000
C.2 EMA ON: Optional Anisotropy Estimation
We also evaluate an optional EMA-based estimation of diagonal precisions (EMA ON), while keeping the prototypes fixed. This variant can be viewed as introducing a light-weight, head-specific geometric reweighting.
Compared to EMA OFF, we observe that EMA can provide marginal gains when $K$ is very small, where prototype coverage is limited. However, once geometric coverage becomes adequate, the performance difference between EMA ON and EMA OFF becomes small and does not change the main trend: routed/oracle performance saturates and FM approaches zero. Accordingly, we treat EMA as an auxiliary variant and use the training-free configuration (EMA OFF) as default.
Table 11: EMA ON. Oracle image-level AUROC across prototype coverage $K$. The last row reports the average over categories.
Category 16 32 64 96 128 196 256 512
Bottle .948 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Cable .736 .853 .925 .936 .948 .975 .979 .987
Capsule .745 .852 .899 .900 .923 .957 .968 .986
Carpet .917 .987 .998 1.00 1.00 1.00 1.00 1.00
Grid .934 .997 .989 .995 1.00 1.00 1.00 1.00
Hazelnut .969 .981 1.00 1.00 1.00 1.00 1.00 1.00
Leather 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
MetalNut .924 .984 .982 .990 .994 .998 1.00 1.00
Pill .869 .904 .949 .931 .940 .958 .959 .969
Screw .522 .498 .644 .728 .720 .855 .887 .949
Tile .957 1.00 1.00 .998 1.00 1.00 1.00 1.00
Toothbrush .781 .994 .969 .994 .983 .975 .967 .986
Transistor .620 .864 .960 .993 .974 .977 .973 .988
Wood .969 1.00 .996 .995 .993 .995 .993 .996
Zipper .884 .964 .973 .970 .983 .986 .988 .992
Mean .852 .925 .952 .962 .964 .978 .981 .990
Table 12: EMA ON. Forgetting measure (FM) across prototype coverage $K$. The last row reports the average over categories.
Category 16 32 64 96 128 196 256 512
Bottle .000 .000 .000 .000 .000 .000 .000 .000
Cable .000 .000 .000 .000 .000 .000 .000 .000
Capsule .000 .000 .000 .000 .000 .000 .000 .000
Carpet .000 .000 .000 .000 .000 .000 .000 .000
Grid .000 .000 .000 .000 .000 .000 .000 .000
Hazelnut .029 .079 .000 .009 .000 .000 .000 .000
Leather .000 .000 .000 .000 .000 .000 .000 .000
MetalNut .000 .000 .000 .000 .000 .000 .000 .000
Pill .273 .000 .000 .000 .000 .000 .000 .000
Screw .012 .000 .000 .000 .000 .000 .000 .000
Tile .000 .000 .000 .000 .000 .000 .000 .000
Toothbrush .000 .000 .000 .000 .000 .000 .000 .000
Transistor .008 .000 .000 .000 .000 .000 .000 .000
Wood .000 .000 .000 .000 .000 .000 .000 .000
Zipper .000 .000 .000 .000 .000 .000 .000 .000
Mean .021 .005 .000 .001 .000 .000 .000 .000
Appendix D Inference Efficiency
To evaluate computational efficiency, we measure inference throughput in frames per second (FPS). To reduce the influence of I/O (e.g., image loading or visualization), we focus on forward-pass latency. Each forward pass includes: (i) feature extraction using a frozen backbone, (ii) task-agnostic routing across heads, and (iii) prototype-based scoring for the routed candidate head.
Inference throughput.
Given $N_{\text{total}}$ test images, forward-pass throughput is
$\mathrm{FPS}_{\text{forward}}=\frac{N_{\text{total}}}{\sum_{i=1}^{N_{\text{total}}}\Delta t_{i}},$ (14)
where $\Delta t_{i}$ denotes the forward-only latency for the $i$-th image. The average per-image latency is
$\mathrm{Latency}\penalty 10000\ (\mathrm{ms})=\frac{1000}{\mathrm{FPS}_{\text{forward}}}.$ (15)
Implementation details.
All measurements are conducted on a single NVIDIA RTX 4090 GPU with input resolution $224\times 224$. We employ a warm-up phase to mitigate CUDA initialization effects. Unless otherwise stated, all results are obtained under task-agnostic inference with routing enabled.
Efficiency–accuracy trade-off.
Table 13 reports throughput under different prototype coverage levels $K$. Latency increases approximately linearly with $K$ due to more prototype comparisons. In our experiments, routing stability saturates at moderate coverage (around $K=64$), while strong detection performance is maintained at $K=196$, offering a practical balance.
Table 13: Forward-pass inference throughput under different prototype coverage $K$. All results are measured on a single RTX 4090 GPU with input resolution $224\times 224$.
$K$ (Prototypes) FPS (images/s) Latency (ms/image)
16 1348.3 0.74
32 1082.2 0.92
64 892.6 1.12
96 718.9 1.39
128 618.2 1.62
196 467.9 2.14
256 370.3 2.70
512 211.5 4.73